# Loop Shaping of Hybrid Motion Control with Contact Transition

### v1

@@ IV-B Evaluated motion control
Two reshaped ’soft’ feedback controllers which allow for a stable and smooth contact transition are designed, cf. section II. The first one, denoted by $C_{v}(s)$, is to enable a purely viscous and well damped repulsive behavior when contacting with environment, and it yields
$u(s)=-\frac{0.0486\,s^{2}+10.78\,s}{(0.0272\,s+4.08)(\omega_{c}^{-1}s+1)}\,x(s).$ (15)
@@
$u(s)=-\frac{0.0486\,s^{2}+10.78\,s}{(0.0272\,s+4.08)(\omega_{c}^{-1}s+1)}\,x(s)+429\,e(s),$ (16)
similar to (15) but differs from this by including also the proportional feedback of the control error, cf. with (14). The latter makes it possible to press on the environment with a force that is either proportional to the control error (in which case the reference value $r(t)$ is to be considered additionally), or with a constant force equal to $U$. For the second case, which is also used in the below experiments, the proportional control part in (16) is extended by saturation, i.e. $\mathrm{sat}_{U}\bigl{[}k_{p}\,e(t)\bigr{]}$. Further we note that both fractions in (15) and (16) use a low-pass filter with a sufficiently high cut-off frequency $\omega_{c}$; otherwise both control transfer functions are improper and could not be implemented, cf. section III.
The resulted magnitude response of the closed-loop disturbance sensitivity function $S(s)$ is compared for all three feedback controllers (here for $r=0$) in Fig. 6.
@@
The experimentally evaluated motion control scenario is shown in Fig. 7 (a). The reference trajectory $r(t)$ with one positive and one negative slope, both implying the same reference velocity magnitude, is tracked by the designed stiff controller (14). On the way back, i.e. at time $t>10$ sec, there is no environmental obstacles and, therefore, no contact with soft objects, cf. with Fig. LABEL:fig:expsetup where a soft object was already placed. Note that at the beginning and especially after the slope segments of trajectory, the control error $|e(t)|$ increases and takes certain time to settle, cf. Fig. 7 (a). This is due to nonlinear friction effects (see [19] for details) which are not explicitly compensated and need to be mitigated by the proportional and integral feedback actions only.
@@
The output of the feedback controller $u(t)$ is depicted in Fig. 7 (b), indicating that it stays in a certain bound. The control reshaping threshold is set to $U=1.3$. Recall that once $|u(t)|>U$, the feedback control changes from its stiff configuration, i.e. (14), to a compliant one, i.e. either (15) or (16) as both controllers $C_{v}$ and $C_{ve}$ were evaluated. Also recall that the reshaping threshold value corresponds to the disturbance force which increases more the stiff motion controller $C_{s}$ is loading the contacting object, cf. section II.
Following to that, the experimental scenarios with a soft environmental object, which is a half of the grape placed before the way back i.e. at time $104<t<$ 10 sec, were evaluated. The final state is exemplified in Fig. LABEL:fig:expsetup. Three control configurations were evaluated. The first one is the stiff control (14) without hybrid reconfiguration to a compliant controller. The second is the stiff control (14) which is reconfigured into the viscous control (15) upon the reshaping threshold $U$. Finally the third is the stiff control (14) which is reconfigured into the viscoelastic control (16) upon the reshaping threshold $U$. Recall that the viscoelastic control (16) is additionally subject to saturation at $|u(t)|=U$ since it contains also feedback proportional to the displacement error $e$.
@@
The measured displacement response $x(t)$ and the feedback control value $u(t)$ are shown in Fig. 8 (a) and (b), respectively. One can recognize that the stiff control $C_{s}$ reaches the back reference position (for $t>10$ sec), thus ploughing the mechanical tool into the soft environment, cf. Fig. LABEL:fig:expsetup (a). Its control value falls below the set threshold and represents the corresponding force required to penetrate into the grape. Quite the opposite, the viscous control $C_{v}$, activated by exceeding the threshold value, provides a slightly repulsive response which can be associated with certain elasticity of the grape surface, cf. Fig. LABEL:fig:expsetup (b). The controller $C_{v}$ maintains the contact position while the control value has a zero mean and almost the same high frequency pattern as the $C_{s}$ control; this is due to the sensor noise and the corresponding output derivative. A slightly differing behavior can be seen in case of the viscoelastic controller $C_{ve}$. Due to a sufficiently large control error $e(t)$ for $t>11$ sec and, at the same time, the used control saturation in $C_{ve}$, the motion system does not penetrate into the grape, but presses it further with the corresponding threshold magnitude. Here we recall that at steady-state, the control value is proportional to the disturbance force coming from the environment.
@@ V Conclusions
In this communication, we provided an easy interpretable (in frequency domain) analysis and design of dynamic reshaping of the stiff (admittance) motion control to a soft (impedance) control upon the contact with constrained and deformable environmental objects. Only the measured output displacement in feedback is used for the overall control design and operation. Hybrid control scheme was established and revealing control experiments with contacting the soft but penetrable objects, i.e. grape, were provided.

### v2

@@ IV-B Evaluated motion control
Two reshaped ’soft’ feedback controllers which allow for a stable and smooth contact transition are designed in accord with section II. The first one, denoted by $C_{v}(s)$, is enabling a purely viscous and well-damped repulsive behavior when contacting with environment, and it yields
$u_{v}(s)=-\frac{0.0486\,s^{2}+10.78\,s}{(0.0272\,s+4.08)(\omega_{c}^{-1}s+1)}\,x(s).$ (15)
@@
$u_{ve}(s)=-\frac{0.0486\,s^{2}+10.78\,s}{(0.0272\,s+4.08)(\omega_{c}^{-1}s+1)}\,x(s)+429\,e(s),$ (16)
which is similar to (15) but differs from this by including also a proportional feedback of the control error, cf. with (14). The latter makes it possible to press on the environment with a force that is either proportional to the control error (in this case the reference value $r(t)$ is to be considered additionally), or with a constant force equal to $U$.
@@
For the second case, which is used in the below experiments, the proportional control part in (16) is extended by saturation, i.e. $\mathrm{sat}_{U}\bigl{[}k_{p}\,e(t)\bigr{]}$. Further we note that both fractions in (15) and (16) use a low-pass filter with a sufficiently high cut-off frequency $\omega_{c}$; otherwise both control transfer functions are improper and could not be implemented, cf. section III.
The resulted magnitude response of the closed-loop disturbance sensitivity functions $S(s)$ are compared for all three feedback controllers (here for $r=0$) in Fig. 6.
The experimentally evaluated motion control scenario is shown in Fig. 7 (a). The reference trajectory $r(t)$ with one positive and one negative slope, both implying the same reference velocity magnitude, is tracked by the designed stiff controller (14). On the way back, at time $t>10$ sec, there is no environmental obstacles and, therefore, no contact with soft objects, cf. with Fig. LABEL:fig:expsetup where a soft object was afterwards placed. Note that at the beginning and especially after the slope segments of trajectory, the control error $|e(t)|$ increases and takes certain time to settle, cf. Fig. 7 (a). This is due to nonlinear friction effects (see [19, 24] for details) which are not explicitly compensated and need to be mitigated by the proportional and integral feedback actions only.
@@
The output of the feedback controller $u(t)$ is depicted in Fig. 7 (b), indicating that it stays in a certain bound. The control reshaping threshold is set to $U=1.3$. Recall that once $|u(t)|>U$, the feedback control changes from its stiff configuration, i.e. (14), to a compliant one, i.e. either (15) or (16) since both controllers $C_{v}$ and $C_{ve}$ were evaluated. Also recall that the reshaping threshold value corresponds to the disturbance force which increases once the stiff motion controller $C_{s}$ is loading the contacting object, cf. section II.
Following to that, the experimental scenarios with a soft environmental object, which is a half of the grape placed before the way back i.e. at time $104<t<$ 10 sec, were evaluated. The final state is exemplified in Fig. LABEL:fig:expsetup. Three control configurations were evaluated. The first one is the stiff control (14) without hybrid reconfiguration to a compliant controller. The second is the stiff control (14) which is reconfigured to the viscous control (15) upon the reshaping threshold $U$. Finally the third is the stiff control (14) which is reconfigured to the viscoelastic control (16) upon the reshaping threshold $U$. Recall that the viscoelastic control (16) is additionally subject to the saturation at $|u(t)|=U$ since it contains also the feedback proportional to $e$.
@@
The measured displacement response $x(t)$ and the feedback control value $u(t)$ are shown in Fig. 8 (a) and (b), respectively. One can recognize that the stiff control $C_{s}$ reaches the back reference position (for $t>10$ sec), thus ploughing the mechanical tool into the soft environment, cf. Fig. LABEL:fig:expsetup (a). Its control value falls below the set threshold and represents the corresponding force required to penetrate into the grape. Quite the opposite, the viscous control $C_{v}$, activated by exceeding the threshold value, provides a slightly repulsive response which can be associated with certain elasticity of the grape surface, cf. Fig. LABEL:fig:expsetup (b). The controller $C_{v}$ maintains the contact position while the control value has a zero mean and almost the same high frequency pattern as the $C_{s}$ control has; this is due to the sensor noise and the corresponding output derivative. A slightly differing behavior can be seen in case of the viscoelastic controller $C_{ve}$. Due to a sufficiently large control error $e(t)$ for $t>11$ sec and, at the same time, the used control saturation in $C_{ve}$, the motion system does not penetrate into the grape, but presses it further with the corresponding threshold magnitude.
The illustrative videos of two of the control experiments reported above, i.e. with the ’stiff’ (PID) control (14) and the ’soft’ viscous control (15), can be found online at https://drive.google.com/file/d/1M9Y3YcCdRDVJ4cOXlq45WOhGyT5Jv4II
https://drive.google.com/file/d/1JrF01yRCI6ZqB6PMSpZ19hkgjczP8zLX respectively.
@@ V Conclusions
In this communication, we provided an easily interpretable (in frequency domain) analysis and design for dynamically transforming stiff (admittance) motion control into soft (impedance) control upon contact with constrained and deformable environmental objects. Only the measured output displacement is used. A hybrid control scheme was established, and insightful control experiments involving contact with soft but penetrable objects, e.g., grapes, were conducted.